OpenStudy (anonymous):

I'm trying to solves the system, but I don't know where to start. Here's the given: (5/x^2) - (2/y^2) + 3 = 0 (3/x^2) - (1/y^2) = 7

OpenStudy (anonymous):

\[\frac{ 5 }{ x ^{2} } - \frac{ 2 }{ y ^{2} } +3 = 0\] \[\frac{ 3 }{ x ^2 } - \frac{ 1 }{ y ^2 } = 7\]

OpenStudy (raden):

times 2 the 2nd equation, what do you get ?

OpenStudy (anonymous):

Oh! \[\frac{ 6 }{ x ^{2} } + \frac{ 2 }{ y ^{2} } = 14\] ?

OpenStudy (raden):

(6/x^2) - (2/y^2) = 14

OpenStudy (anonymous):

I have it now! Thank you!

OpenStudy (anonymous):

Put 1/x^2=a and 1/y^2=b...the equations will turn into linear equation in two variables which you can solve simuntaneously